40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 25 02:03:59 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts 
    Reply  Fri Jul 28 15:36:32 2017, gautam, Update, Optical Levers, Optical lever tuning thoughts errSig.pdfQPD_simulink.pngfeedbackTopology.pdf
       Reply  Tue Aug 1 16:05:01 2017, gautam, Update, Optical Levers, Optical lever tuning - cost function construction loopPerformance.pdfOplevLoop.pngOL_errSigs.pdfDAC_saturation.pdf
          Reply  Thu Nov 16 13:57:01 2017, gautam, Update, Optical Levers, Optical lever noise VertexOLnoise.pdfITMYvsSRM.pdf
             Reply  Thu Nov 16 15:43:01 2017, rana, Update, Optical Levers, Optical lever noise 
                Reply  Tue Nov 21 11:37:29 2017, gautam, Update, Optical Levers, BS OL calibration updated OL_calib_BS_PERROR.pdfOL_calib_BS_YERROR.pdfVertexOLnoise_updated.pdf
                   Reply  Tue Nov 21 16:28:23 2017, gautam, Update, Optical Levers, BS OL calibration updated BS_oplev_sweep.png
                      Reply  Tue Nov 21 23:04:12 2017, gautam, Update, Optical Levers, Oplev "noise budget" vertexOL_noises.pdf
                         Reply  Wed Nov 22 05:41:32 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                            Reply  Wed Nov 22 15:29:23 2017, gautam, Update, Optical Levers, Oplev "noise budget" OplevTest.jpg
                               Reply  Wed Nov 22 16:40:00 2017, Koji, Update, Optical Levers, Oplev "noise budget" 
                               Reply  Wed Nov 22 19:20:01 2017, rana, Update, Optical Levers, Oplev "noise budget" 
                                  Reply  Wed Nov 22 23:56:14 2017, gautam, Update, Optical Levers, Oplev "noise budget" 
                                     Reply  Thu Nov 23 18:03:52 2017, gautam, Update, Optical Levers, Oplev "noise budget" ITMY_P_noise.pdfITMY_Y_noise.pdf
Message ID: 13437     Entry time: Tue Nov 21 11:37:29 2017     In reply to: 13433     Reply to this: 13439
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: BS OL calibration updated 

I calibrated the BS oplev PIT and YAW error signals as follows:

  1. Locked X-arm, ran dither alignment servos to maximize transmission.
  2. Applied an offset to the ASC PIT/YAW filter banks. Set the ramp time to something long, I used 60 seconds.
  3. Monitored the X arm transmission while the offset was being ramped, and also the oplev error signal with its current calibration factor.
  4. Fit the data, oplev error signal vs arm transmission, with a gaussian, and extracted the scaling factor (i.e. the number which the current Oplev error signals have to be multiplied by for the error signal to correspond to urad of angular misalignment as per the overlap of the beam axis to the cavity axis.
  5. Fits are shown in Attachment #1 and #2.
  6. I haven't done any error analysis yet, but the open loop OL spectra for the BS now line up better with the other optics, see Attachment #3 (although their calibration factors may need to be updated as well...). Need to double check against OSEM readout during the sweep.
  7. New numbers have been SDF-ed.

The numbers are:

BS Pitch     15  /  130    (old/new)     urad/counts

BS Yaw       14  /  170    (old/new)     urad/counts

Quote:

I bet the calibration is out of date; probably we replaced the OL laser for the BS and didn't fix the cal numbers. You can use the fringe contrast of the simple Michelson to calibrate the OLs for the ITMs and BS.

 

Attachment 1: OL_calib_BS_PERROR.pdf  40 kB  | Hide | Hide all
OL_calib_BS_PERROR.pdf
Attachment 2: OL_calib_BS_YERROR.pdf  44 kB  | Hide | Hide all
OL_calib_BS_YERROR.pdf
Attachment 3: VertexOLnoise_updated.pdf  81 kB  | Hide | Hide all
VertexOLnoise_updated.pdf
ELOG V3.1.3-